40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 10 01:25:28 2020, gautam, Update, LSC, Locking notes 
    Reply  Fri Jul 10 15:26:28 2020, gautam, Update, LSC, MC2 coils need DC balancing? MC2_balancing.pngASDC.png
       Reply  Sat Jul 11 18:24:30 2020, Koji, Update, LSC, MC2 coils need DC balancing? 
       Reply  Sun Jul 12 02:42:01 2020, gautam, Update, LSC, Locking (on rossa) ASCsteps.png
       Reply  Mon Jul 13 11:36:08 2020, rana, Update, LSC, MC2 coils need DC balancing? 
Message ID: 15466     Entry time: Fri Jul 10 01:25:28 2020     Reply to this: 15468
Author: gautam 
Type: Update 
Category: LSC 
Subject: Locking notes 

More tomorrow, but I tried the following tonight:

  1. Dither alignment for PRC / MICH seems to work when the PRFPMI is locked. Unfortunately, the correct settings for the arm cavity dither alignment loops continue to elude me.
  2. I tried some arm ASC loop characterization by stepping the error points of these loops - I saw some weird cross coupling between the loops that needs investigation.
  3. I'm unable to turn an integrator on for the "Common YAW" QPD loop - unclear why this is, but every time I attempt to engage said integrator, the lock is immediately blown. Needs some investigation of the signals.
  4. With the PRC / MICH angular DoFs aligned with the dither alignemnt, and the arm ASC loops hand tuned, I was able to get the darkest dark port I've seen. In terms of ASDC counts, it was ~ 200, which after undoing all the digital gains etc corresponds to ~100 uW of light. I think we can get a rough estimate of the contrast defect by accounting for (i) T_SRM, (ii) OMC pickoff fraction (iii) other losses between the BS dark port and the AP table (iv) 50/50 BS between AS55 and AS110 PD (the ASDC signal is derived from the former) and (v) the throughput of the 55 MHz sideband to the dark port, although there are many uncertainties. 
ELOG V3.1.3-